# Lecture Notes on Spectral Independence and Bases of a Matroid: Local-to-Global and Trickle-Down from a Markov Chain Perspective

### v3

@@ 8.5 Application: Reliability
An interesting application is to compute reliability. Let $X\subseteq[n]$ be a random set where every element is included independently with probability $p$. The reliability is defined as the $\mathbb{P}[X\supseteq\text{basis}]$, and the unreliability is defined as $\mathbb{P}[X\not\supseteq\text{basis}]$.
@@
Theorem 6 yields an $\mathsf{FPRAS}$ for the reliability of a matroid. In particular, rapid mixing of the bases-exchange walk yields an efficient sampling algorithm for generating a basis of a matroid (almost) uniformly at random; the classical result of [JVV86] then yields an $\mathsf{FPRAS}$ for computing the number of bases.
Consider a matroid $\mathcal{M}$ for which we would like to compute the reliability, which is the probability that the random subset $X$ contains a basis of $\mathcal{M}$. Let $\mathcal{M}^{*}$ denote the dual of $\mathcal{M}$, see 8.3; note $\mathcal{M}^{*}$ is a matroid. Note,
$\mathrm{Prob}\left(X\supseteq\text{basis of $\mathcal{M}$}\right)=\mathrm{Prob}\left(\overline{X}\subseteq\text{basis of $\mathcal{M}^{*}$}\right)=\sum_{k}(1-p)^{k}p^{n-k}\text{\#\{indep set of $\mathcal{M}^{*}$ of size $k$\}},$
@@ 9 Trickle-Down Theorem
Here we present Oppenheim’s Trickle-Down Theorem [Opp18] which is the new tool needed to prove Theorem 6. We present the theorem in the general context of simplicial complexes; this enables one to apply it both in the context of spin systems and for bases of a matroid.
@@ 9.2 Trickle-Down Statement
Theorem 7 ([Opp18]).
Consider $S\in\Omega$. Let $i=|S|$ and assume $20\leq i<r-$ 2. Let $\gamma_{i}=\gamma(Q_{S})$. Assume that $Q_{S}$ is irreducible and hence $\gamma_{i}>0$. Suppose there exists $\gamma_{i+1}>0$ where for all $Z\in\Omega$ such that $S\subset Z$ and $|Z|=|S|+1$,
@@
10 Rapid Mixing of Bases-Exchange Walk: Proof of Theorem 6
The proof of Theorem 6 uses the Trickle Down Theorem (Theorem 7), which bounds the spectral gap of the local walks, and the Random Walk Theorem (Theorem 4), which is the local to global theorem.
Our goal is to bound the spectral gap of the bases-exchange walk using the Random Walk Theorem. To that end we need to bound the spectral gap of the local walk (this is the up-down walk $\mathrm{P}_{1}^{\wedge}$) for every pinning (or link in the terminology of simplicial complexes). The pinnings are defined for every independent set $S$. In particular, for an independent set $S$ the pinning is defined as $\{B\setminus S:S\subseteq B,B\in\mathcal{I}\}$, which is a contraction of the matroid. 8.5 establishes that the contraction of a matroid is also a matroid.
These local walks for a particular pinning are weighted as these are the projections in the original simplicial complex. To handle these weighted local walks we utilize the Trickle Down Theorem. The Trickle Down Theorem says that if we have an “optimal” bound on the spectral gaps for all local walks on $(r-2)$-dimensional pinning/link then we obtain an optimal bound on the spectral gap for the local walks for all pinnings/links, and then we can apply the Random Walk Theorem. This greatly simplifies matters because the local walks on $(r-2)$-dimensional pinnings/links are unweighted, and hence they simply correspond to (unweighted) rank 2 matroids.
@@ 10 Rapid Mixing of Bases-Exchange Walk: Proof of Theorem 6
The adjacency matrix $A$ of a complete multipartite graph is the all ones matrix (that has only one non-zero eigenvalue) minus a block diagonal matrix, so it has the second eigenvalue $\leq 0$. (Adding the isolated vertices doesn’t change the non-zero part of the spectrum.)
The local walk $Q_{\mathcal{M}}$ is equivalent to the random walk on the graph $G_{\mathcal{M}}$. The transition matrix of the random walk on $G_{\mathcal{M}}$ is the matrix $D^{-1}A$ where $D$ is the diagonal matrix with the degrees. Note, $D^{-1}A$ is similar to $D^{-1/2}AD^{-1/2}$. Sylvester’s law of inertia says that for matrices $A$ and $B$, the signature of $A$ and $BAB^{T}$ are the same, where the signature of a matrix refers to the number of positive and negative eigenvalues. Hence, from Sylvester’s law of inertia (which states that a change of basis does not change the number of eigenvalues of a given sign) we have that $\lambda_{2}(A)\leq 0\iff\lambda_{2}(D^{-1/2}AD^{-1/2})\leq 0$. This completes the proof. ∎
@@
This lemma follows from Lemma 10.1 and Corollary 9.1 (since the $(r-2)$-links are rank $2$ matroids). ∎
@@
Proof of Theorem 6.
The random walk theorem Theorem 4 implies that the spectral gap $\gamma$ for the transition matrix $P$ of the bases exchange walk satisfies:
@@
11 Proof of the Trickle-Down Theorem: Proof of Theorem 7
In this section, we complete the proof of the Trickle-Down Theorem Theorem 7; this is the only remaining task to complete the proof of fast mixing of the bases-exchange walk (Theorem 6).
@@ 11.1 Proofs of Technical Lemmas
Recall $i=|S|$, and $\mu$ is the uniform distribution over subsets $S\subset E$ which are bases of the matroid. Let us begin with two basic facts.
$\pi_{S}(a)=\pi_{S,1}(a)=\frac{\mu(S\cup a)}{(n-i)\mu(S)}=\frac{\pi_{i+1}(S\cup a){n\choose i+1}}{(n-i)\pi_{i}(S){n\choose i}}=\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}$ (56)
@@
We can now proceed with the proof of Eq. 54.
@@
We can establish Eq. 53 in a similar manner.
@@
$=\sum_{a\in E\setminus S}\pi_{S}(a)\mathcal{E}_{Q_{S\cup a}}(f).$
We now prove Lemma 11.2, which will complete the proof of the Trickle-Down Theorem (Theorem 7).

### v4

@@ 7.5 Application: Reliability
An interesting application is to compute reliability. Let $X\subseteq[n]$ be a random set where every element is included independently with probability $p$. The reliability is defined as the $\mathrm{Pr}(X\supseteq\text{basis})$, and the unreliability is defined as $\mathrm{Pr}(X\not\supseteq\text{basis})$.
@@
Theorem 1.15 yields an $\mathsf{FPRAS}$ for the reliability of a matroid. In particular, rapid mixing of the bases-exchange walk yields an efficient sampling algorithm for generating a basis of a matroid (almost) uniformly at random; the classical result of [JVV86] then yields an $\mathsf{FPRAS}$ for computing the number of bases.
Consider a matroid $\mathcal{M}$ for which we would like to compute the reliability, which is the probability that the random subset $X$ contains a basis of $\mathcal{M}$. Let $\mathcal{M}^{*}$ denote the dual of $\mathcal{M}$, see 7.3; note $\mathcal{M}^{*}$ is a matroid. Note,
[…]
@@ 8 Trickle-Down Theorem
Here we present Oppenheim’s Trickle-Down Theorem [Opp18] which is the new tool needed to prove Theorem 1.15. We present the theorem in the general context of simplicial complexes; this enables one to apply it both in the context of spin systems and for bases of a matroid.
@@ 8.2 Trickle-Down Statement
Theorem 8.1 ([Opp18]).
Consider $S\in\Omega$. Let $i=|S|$ and assume $0\leq i<r-2$. Let $\gamma_{i}=\gamma(Q_{S})$. Assume that $Q_{S}$ is irreducible and hence $\gamma_{i}>0$. Suppose there exists $\gamma_{i+1}>0$ where for all $Z\in\Omega$ such that $S\subset Z$ and $|Z|=|S|+1$,
@@
8.3 Rapid Mixing of Bases-Exchange Walk: Proof of Theorem 1.15
The proof of Theorem 1.15 uses the Trickle-down Theorem (Theorem 8.1), which bounds the spectral gap of the local walks, and the Random Walk Theorem (Theorem 4.1), which is the local to global theorem.
Our goal is to bound the spectral gap of the bases-exchange walk using the Random Walk Theorem. To that end we need to bound the spectral gap of the local walk (this is the up-down walk $\mathrm{P}_{1}^{\wedge}$) for every pinning (or link in the terminology of simplicial complexes). The pinnings are defined for every independent set $S$. In particular, for an independent set $S$ the pinning is defined as $\{B\setminus S:S\subseteq B,B\in\mathcal{I}\}$, which is a contraction of the matroid. 7.5 establishes that the contraction of a matroid is also a matroid.
These local walks for a particular pinning are weighted as these are the projections in the original simplicial complex. To handle these weighted local walks we utilize the Trickle-down Theorem. The Trickle-down Theorem says that if we have an “optimal” bound on the spectral gaps for all local walks on (r2)(r-2)-dimensional pinning/link then we obtain an optimal bound on the spectral gap for the local walks for all pinnings/links, and then we can apply the Random Walk Theorem. This greatly simplifies matters because the local walks on (r2)(r-2)-dimensional pinnings/links are unweighted, and hence they simply correspond to (unweighted) rank 2 matroids.
@@
The adjacency matrix $A$ of a complete multipartite graph is the all ones matrix (that has only one non-zero eigenvalue) minus a block diagonal matrix, so it has the second eigenvalue $\leq 0$. (Adding the isolated vertices does not change the non-zero part of the spectrum.)
The local walk $Q_{\mathcal{M}}$ is equivalent to the random walk on the graph $G_{\mathcal{M}}$. The transition matrix of the random walk on $G_{\mathcal{M}}$ is the matrix $D^{-1}A$ where $D$ is the diagonal matrix with the degrees. Note, $D^{-1}A$ is similar to $D^{-1/2}AD^{-1/2}$. Hence, by Sylvester’s law of inertia, see Lemma 2.1 (which states that a change of basis does not change the number of eigenvalues of a given sign) we have that $\lambda_{2}(A)\leq 0\iff\lambda_{2}(D^{-1/2}AD^{-1/2})\leq 0$. This completes the proof. ∎
@@
This lemma follows from Lemma 8.3 and Corollary 8.2 (since the (r2)(r-2)-links are rank $2$ matroids). ∎
@@
Proof of Theorem 1.15.
The random walk theorem Theorem 4.1 implies that the spectral gap $\gamma$ for the transition matrix $P$ of the bases-exchange walk satisfies:
@@
8.4 Connections to Spectral Independence and Log-Concavity
Recall, the definition of the modified influence matrix $\widetilde{\Psi}$ in Definition 2.7, which is slightly different than the influence matrix $\Psi$ in Definition 1.1; the former $\widetilde{\Psi}$ leads to the notion of weak spectral independence (see Definition 2.9) and the latter $\Psi$ leads to spectral independence (see Definition 1.2). Both weak spectral independence and spectral independence upper bound the maximum eigenvalue $\lambda_{\max}$ of the corresponding matrix but do so for all valid pinnings (which corresponds to all induced subgraphs in the case of the hard-core model on independent sets).
We first observe that the local walks having an optimal spectral gap is equivalent to optimal spectral independence.
Remark 8.5.
The following are equivalent by Lemma 4.2:
  1. 1.
    For every pinning $S$, the local walk $Q_{S}$ satisfies $\gamma(Q_{S})\geq 1$.
  2. 2.
    0-spectral independence holds.
Recall, 0-spectral independence implies 0-weak spectral independence. Now we relate 0-weak spectral independence with log-concavity.
Given a distribution $\mu$ on $\{0,1\}^{n}$ consider the following generating polynomial
$f(x_{1},\dots,x_{n})=\sum_{\sigma\in\{0,1\}^{n}}\mu(\sigma)\prod_{i=1}^{n}x_{i}^{\sigma_{i}}.$
Definition 8.6 (log-concavity at a point).
We say that $f$ is log-concave at a point $x$ if $(\nabla^{2}\log f)(x)\preceq 0$.
As noted above, 0-weak spectral independence is defined as $\lambda_{\max}(\widetilde{\Psi_{\tau}})\leq 1$ for all pinnings $\tau$. The following lemma shows that $\lambda_{\max}(\widetilde{\Psi})\leq~1$ is equivalent to log-concavity of $f$.
Theorem 8.7.
Log-concavity of $f$ at the point $(1,\dots,1)$ is equivalent to the influence matrix $\widetilde{\Psi}$ having $\lambda_{\max}(\widetilde{\Psi})\leq~1$.
Proof.
For $1\leq i\neq j\leq n$ we have
$x_{i}x_{j}(\nabla^{2}\log f(1,\dots,1))_{ij}$ $=x_{i}x_{j}\left(\frac{(\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}f)}{f}(1,\dots,1)-\frac{(\frac{\partial}{\partial x_{i}}f)(\frac{\partial}{\partial x_{i}}f)}{f^{2}}(1,\dots,1)\right)$ (84)
$=\mathrm{Pr}_{\sigma\sim\mu}(\sigma(i)=\sigma(j)=1)-\mathrm{Pr}_{\sigma\sim\mu}(\sigma(i)=1)\mathrm{Pr}_{\sigma\sim\mu}(\sigma(j)=1),$ (85)
and for $1\leq i\leq n$ we have
$x_{i}^{2}(\nabla^{2}\log f(1,\dots,1))_{ii}$ $=-x_{i}^{2}\left(\frac{\frac{\partial}{\partial x_{i}}f}{f}(1,\dots,1)\right)^{2}=-\mathrm{Pr}_{\sigma\sim\mu}(\sigma(i)=1)^{2}.$ (86)
Recall from Eq. 4 the definition of the covariance matrix:
$\mathrm{Cov}_{\mu}(i,j)$ $=\mathrm{Pr}_{\sigma\sim\mu}[\sigma(i)=\sigma(j)=1]-\mathrm{Pr}_{\sigma\sim\mu}[\sigma(i)=1]\cdot\mathrm{Pr}_{\sigma\sim\mu}[\sigma(j)=1].$
Let $\widetilde{D}={\mathrm{diag}}(\mathbb{E}(\mu))$ be the diagonal matrix with entries $\widetilde{D}(i,i)=\mu(\sigma(i)=1)$. Hence $(\nabla^{2}f)(1,\dots,1)\preceq 0$ is equivalent to
$\mathrm{Cov}_{\mu}-\widetilde{D}=\mathrm{diag}(x_{i})(\nabla^{2}f(1,\dots,1))\mathrm{diag}(x_{i})\preceq 0,$ (87)
where in the first equality we used Eq. 84 and (86).
Recall from Eq. 13 that $\lambda_{\max}(\widetilde{\Psi})\leq 1$ is equivalent to $\mathrm{Cov}_{\mu}\preceq\widetilde{D}$. Therefore, Eq. 87 establishes the condition Eq. 12 in the definition of 0-weak spectral independence, see Definition 2.9. ∎
We refer interested readers to the works [ALOV19, BH20] for more discussion on (strongly) log-concave polynomials.
8.5 Proof of the Trickle-Down Theorem: Proof of Theorem 8.1
In this section, we complete the proof of the Trickle-Down Theorem (Theorem 8.1); this is the only remaining task to complete the proof of fast mixing of the bases-exchange walk (Theorem 1.15).
Remark 8.8.
For an alternative proof of the Trickle-Down Theorem (Theorem 8.1) in the language of covariance matrices (see Eq. 10), see [Koe24, AKV24]. We highlight that [AKV24] established a trickle-down theorem for general localization schemes, which can be applied to the analysis of the field dynamics for the hardcore model; see Section 11 for more details.
@@ 8.6 Proofs of Technical Lemmas
Recall $i=|S|$, and $\mu$ is the uniform distribution over subsets $S\subset E$ which are bases of the matroid. Let us begin with two basic facts. Observe that
$\pi_{S}(a)=\pi_{S,1}(a)=\frac{\mu(S\cup a)}{(n-i)\mu(S)}=\frac{\pi_{i+1}(S\cup a)\binom{n}{i+1}}{(n-i)\pi_{i}(S)\binom{n}{i}}=\frac{\pi_{i+1}(S\cup a)}{(i+1)\pi_{i}(S)}.$ (91)
@@
We can now proceed with the proof of Eq. 89:
@@
We can establish Eq. 88 in a similar manner:
@@
$=\sum_{a\in E\setminus S}\pi_{S}(a)\mathcal{E}_{Q_{S\cup a}}(f).\qed$
We now prove Lemma 8.10, which will complete the proof of the Trickle-Down Theorem (Theorem 8.1).